OpenStudy (anonymous):

Determine an exact value for each expression 1) sin pie/3 tan pie/6 DIVIDED BY cos pie/4

OpenStudy (anonymous):



OpenStudy (ranga):

BTW, It is usually written as pi not pie :)

OpenStudy (jziggy):

\[(\sin(\pi/3) * \tan(\pi/6))/\cos(\pi/4)\] (sqrt(3)/2 * 1/sqrt(3))/sqrt(2)/2 (1/2)/(sqrt(2)/2) 1/sqrt(2)

OpenStudy (anonymous):

answer is sqroot 2/ 2

OpenStudy (jziggy):

if you multiply 1/sqrt(2) by sqrt(2)/sqrt(2) which is also equal to 1 you get sqrt(2)/2

OpenStudy (jziggy):

Some teachers prefer there to not be any radicals in the denominator but in reality 1/sqrt(2) and sqrt(2)/2 are the same

OpenStudy (anonymous):

wow thanks :)